A centrifugal pump operating under steady flow conditions delivers 2, 270 kgm/min of water

from an initial pressure of 82, 740 Pa to a final pressure of 275, 800 Pa. The diameter of the inlet

pipe to the pump is 15.24 cm and the diameter of the discharge pipe is 10.16 cm. Determine the

work. Is it done on or done by the system?



1
Expert's answer
2021-11-19T10:49:32-0500

"H=\frac{p_2-p_1}{\rho g}+\frac{v_2^2-v_1^2}{2g},"

"\dot m=\rho A_1v_1=\rho A_2v_2," "\implies"

"v_1=2.074~\frac ms,~v_2=4.667~\frac ms,"

"H=20.6~m."